Grade 12Physical Chemistry

Calculate the de Broglie wavelength of an electron that has been accelerated by a potential difference of 500 V.

Vatsalya Choudhary
8 Years agoGrade 12

1 Answer

ashish
8 Years ago
Lambda=12.3/√V angstromWhere V=accelerating potentialTherefore ans= 12.3/√500. = 0.55 angstrom i.e. 0.55 × 10^(-10) metreDONT FORGET TO UPVOTE😀